OpenStudy (anonymous):

A car rental company's standard charge includes an initial fee plus an additional fee for each mile driven. The standard charge S (in dollars) is given by the function =S+17.75+0.60M, where M is the number of miles driven.The company also offers an option to insure the car against damage. The insurance charge I (in dollars) is given by the function =I=3.70+0.25m. Let C be the total charge (in dollars) for a rental that includes insurance. Write an equation relating C to M . Simplify your answer as much as possible.

OpenStudy (anonymous):

@Directrix plz help

Directrix (directrix):

Is this supposed to be S = rather than = S >> =S+17.75+0.60M, This "=S+17.75+0.60M," is not an equation.

Directrix (directrix):

Same question here: =I=3.70+0.25m. @Missiey

OpenStudy (anonymous):

yes stupid computer

Directrix (directrix):

S = 17.75 +.6*M where M I = 3.70 + .25*M C = S + I C = 17.75 + .6*M + 3.70 + .25*M @Missiey Add 17.75 to 3.70 Then, add .6M to .25 M Post what you get, okay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c=21.45+0.85m

Directrix (directrix):

Usually, the variable is written first (doesn't have to be). C = .85*M + 21.45 That would be the cost of the car rental and insurance based on miles the car is driven.

OpenStudy (anonymous):

thank you for explaining

Directrix (directrix):

You are welcome.